OpenStudy (anonymous):

what should you substitute for x in the second equation order to solve the system by the substitution method? write your answer as a polynomial in descending order of terms. 1/2x+4y=6 x-3y=9

OpenStudy (anonymous):

\[x/2 = 6 - 4y\] x = 12 - 8y should be substituted in second equation...

OpenStudy (anonymous):

is it 1/(2x) or x/2